# Masks and Boxes: Combining the Best of Both Worlds for Multi-Object Tracking

### v2

@@ Appendix B State-of-the-art comparison with offline and transformer-based methods
MOTR [37] 57.8 73.4 68.6
MeMOTR [13] 58.8 72.8 71.5
MOTRv2 [39] 62.0 78.6 75.0
MOTIP [14] 59.2 75.5 71.2
@@
SUSHI [5] 66.5 81.1 83.1
@@
ByteTrack [38] 63.1 80.3 77.3
StrongSORT++ [12] 64.4 79.6 79.5
OC-SORT [4] 63.2 78.0 77.5
Deep OC-SORT [27] 64.9 79.4 80.6
Hybrid-SORT [36] 64.0 79.9 78.7
C-BIoU [35] * 62.4 79.5 77.1
McByte (ours) 64.2 80.2 79.4
@@
Method HOTA MOTA IDF1
@@
MOTR [37] 54.2 79.7 51.5
MeMOTR [13] 63.4 85.4 65.5
MOTRv2 [39] 73.4 92.1 76.0
MOTIP [14] 67.5 90.3 72.2
@@
SUSHI [5] 63.3 88.7 63.4
@@
ByteTrack [38] 47.7 89.6 53.9
OC-SORT [4] 55.1 92.2 54.9
Deep OC-SORT [27] 61.3 92.3 61.5
C-BIoU [35] * 45.8 88.4 52.0
StrongSORT++ [12] 55.6 91.1 55.2
Hybrid-SORT [36] 65.7 91.8 67.4
McByte (ours) 67.1 92.9 68.1
@@
There are MOT methods outside the tracking-by-detection domain that perform better than ours on some benchmarks, but usually they are not directly comparable, because they make certain hypotheses, e.g. global optimization on the whole video. At the same time, these methods perform visibly worse on other benchmarks as we discuss below. On the contrary, we stress that McByte performs quite well on all the discussed benchmarks (Secs. 4.3 and 4.4 of the main paper). McByte is a tracking-by-detection approach, which is the main focus of our work. For an additional reference, though, we also list performance of the transformer-based and global optimization methods.
@@
Tab. 10 shows extended state-of-the art comparison on MOT17 [28] test set. Transformer-based methods perform visibly lower than the tracking-by-detection methods (including ours) as they struggle with the subjects frequently entering and leaving the scene. In contrast, SUSHI [5], which is a powerful global optimization approach, reaches highly satisfying performance. However, it accesses all the video frames at same time while processing detections and associating the tracklets, which makes it impossible to run in online settings. This makes the performance not directly comparable and therefore, we do not include it in the main paper.
@@
Tab. 11 presents extended state-of-the-art comparison on DanceTrack [33] test set. As in this dataset the subjects remain mostly at the scene, the transformer-based methods performance is more satisfying. The performance of transformer-based methods can be both higher [39, 14] or lower [37, 13] compared to the the tracking-by-detection methods. For similar reasons, the global optimization method, SUSHI [5] can also perform higher than the other tracking-by-detection methods on this dataset, or lower, e.g. when compared to our method. As in the case of the MOT17 benchmark, the performance is not directly comparable and thus not included in the main paper.
Assumed MOT17 [28] sequence correspondence
MOT17 [28] test sequence 01 03 06 07 08 12 14
ByteTrack [38] thresholds 0.65 0.6 0.65 0.6 0.6 0.7 0.67
Table 12: Different high confidence detection thresholds applied by ByteTrack [38] and most of its successors per test sequences of the MOT17 [28] dataset.
Assumed MOT17 [28] sequence correspondence
MOT17 test sequence 01 03 06 07 08 12 14
MOT17 train/val sequence 02 04 05 10 09 11 13
Table 13: Our assumed correspondence between training/validation and test sequences of the MOT17 [28] dataset based on the provided descriptions and their characteristics.
HOTA per sequence
thresholds overall 02 04 05 09 10 11 13
0.5 67.4 48.6 78.8 61.4 66.6 54.8 62.0 64.4
0.6 68.4 50.8 79.6 61.6 67.6 54.4 63.2 65.5
0.7 68.3 50.7 79.7 58.9 66.4 52.2 66.8 62.0
from ByteTrack [38] 68.0 46.1 79.6 61.1 67.6 54.4 66.8 63.5
Table 14: Ablation study: HOTA comparison of the baseline [38] on the MOT17 [28] validation set per sequence and on all sequences (overall) with varying high confidence detection threshold parameters. The highest scores per column are in bold and marked in red, whereas the lowest scores are in bold and marked in blue.
MOTA per sequence
thresholds overall 02 04 05 09 10 11 13
0.5 77.2 54.9 88.7 79.1 86.5 71.2 66.6 74.8
0.6 78.2 58.0 89.0 78.3 86.4 69.6 71.2 77.0
0.7 77.3 58.4 89.9 75.6 84.6 58.7 74.8 73.4
from ByteTrack [38] 78.3 57.1 89.0 78.3 86.3 69.6 74.8 77.0
Table 15: Ablation study: MOTA comparison of the baseline [38] on the MOT17 [28] validation set per sequence and on all sequences (overall) with varying high confidence detection threshold parameters. The highest scores per column are in bold and marked in red, whereas the lowest scores are in bold and marked in blue.
IDF1 per sequence
thresholds overall 02 04 05 09 10 11 13
0.5 78.9 58.7 90.7 76.0 78.6 69.2 69.5 83.0
0.6 80.2 61.4 91.6 76.6 79.6 68.5 71.6 84.3
0.7 80.1 60.6 91.7 73.8 78.7 66.7 76.0 79.4
from ByteTrack [38] 79.8 57.1 91.6 76.0 79.6 68.5 76.0 81.7
Table 16: Ablation study: IDF1 comparison of the baseline [38] on the MOT17 [28] validation set per sequence and on all sequences (overall) with varying high confidence detection threshold parameters. The highest scores per column are in bold and marked in red, whereas the lowest scores are in bold and marked in blue.
@@
Figure 5: An example of detection quality difference. It can be seen that considerably more bounding boxes are missed in case of public detections, which negatively impacts the MOT performance. Input image data from [28], sequence MOT17-04, last frame (1050). Best seen in color.
Appendix C More details on our C-BIoU implementation and performance
The C-BIoU [35] paper does not provide a public implementation or sufficient details for code reproduction. We implement C-BIoU based on the available information and report the best results on the datasets evaluated by the original algorithm. These results, marked with an asterisk, are shown in the main paper (Tabs. 3, 4 and 5) and in Tabs. 10 and 11 of this supplementary material.
Based on the descriptions in the C-BIoU work [35], especially regarding tracklet management and tracklet-detection association, the core of the method strongly resembles ByteTrack [38]. Thus, we extend ByteTrack by incorporating cascaded buffered intersection over union for the association process. Where C-BIoU does not specify parameters, we default to those used in ByteTrack. For the MOT17 test set, we do not tune detection confidence parameters per sequence (as explained in Sec. 2.2 of the main paper) and apply a fixed 0.6 threshold for all sequences and datasets. Regarding the two buffering scales in C-BIoU, we use the values 0.3 and 0.4, which were reported to result in the best HOTA on the DanceTrack validation set in the C-BIoU paper. Although we also test the values 0.3 and 0.5 (as mentioned in the framework figure from C-BIoU), the 0.3 and 0.4 combination performs slightly better in our experiments.
Additionally, following C-BIoU’s supplementary material, we adopt the most optimal max_age values per dataset. The max_age parameter defines the maximum lifespan (in frames) of a tracklet that has not been matched to a detection before it is terminated and removed. For DanceTrack [33], we set max_age to 100, and for SoccerNet-tracking 2022 [9], we set it to 60. In the case of MOT17 [28], where this value is not directly specified, we test three different max_age values: 20, 60, and 100 (as listed in C-BIoU’s supplementary material). We select the value that results in the highest performance on the MOT17 validation set, which is 60, and then use this value for testing on the MOT17 test set to report the final results.
The results are presented in Tabs. 3, 4 and 5 of the main paper. Compared to the baseline, ByteTrack [38], the improvement is modest (Tab. 5) or shows slight degradation (Tabs. 3 and 4). This could be due to tuning issues, as many parameters were not specified in the C-BIoU paper. Specifically, with buffered intersection over union (IoU), the association search space expands, leading to more potential associations and lower IoU-based distances. However, the matching thresholds for the linear association steps (using the Hungarian Algorithm [21]), which filter out unlikely matches, are not provided. Default thresholds from ByteTrack may allow more matches, introducing unresolved ambiguities and potentially degrading performance.
Appendix D Parameter tuning an its impacts
When parameters in tracking-by-detection algorithms [38, 12, 4, 27, 36] are varied, the impact on results can be considerable. We demonstrate this using the MOT17 dataset [28]. Tab. 16 shows different detection confidence threshold values used per MOT17 test sequence by the baseline [38]. Successor methods [12, 4, 27, 36] use the same code base, including these hard-coded parameters (see Sec. 2.2 in the main paper).
We apply both these hard-coded values and a single threshold for all sequences (0.5, 0.6, and 0.7) to ByteTrack [38], evaluating it on the MOT17 validation set. The validation set is made up of the second half of each sequence from the training set. While parameters are typically tuned for the test set in the community, given the test server’s submission limits [10], we match training/validation sequences with their corresponding test sequences. This matching is based on the descriptions and characteristics provided in [28], and we list it inTab. 16.
Tabs. 16, 16 and 16 present ByteTrack’s performance with these parameter sets for HOTA, MOTA, and IDF1 scores. As demonstrated, even small changes in parameter values can cause noticeable variations in results, both overall and for individual sequences. We highlight the best scores in red and the lowest in blue.
Tracking results can be highly sensitive to hyperparameters. With parameters tuned per test sequence, tracking-by-detection algorithms [38, 12, 4, 27, 36] achieve high performance, particularly on MOT17. However, tuning parameters per sequence would be laborious and impractical for datasets with many test sequences, like DanceTrack [33] (35 sequences), SoccerNet-tracking 2022 [9] (49 sequences), and KITTI-tracking [16] (29 sequences).
In McByte, we maintain consistent parameters across a dataset, only adjusting them when detection quality differs significantly, i.e., when using PermaTr [34] detections in KITTI-tracking (which still involves using one value for all sequences of a dataset). Despite not having its parameters tuned per sequence, McByte improves upon the baseline [38] and performs comparably to other tracking-by-detection methods that tune parameters for each sequence on the MOT17 dataset [28].
Appendix E McByte and ByteTrack on public detections
@@
We also evaluate McByte and the baseline [38] on the MOT17 [28] validation set using the public detections provided by the dataset authors. The results are shown in Tab. 17, using the FRCNN model variant [30]. There are significant performance differences for both the baseline and McByte when using public detections, which are of lower quality compared to the private YOLOX [15] detections, specifically tuned on MOT17 by the baseline [38]. Public detections tend to miss more objects, leading to many false negatives.
@@
An example comparing the quality of private and public detections is illustrated in Fig. 5. Since tracklets and their masks can only be initiated when a bounding box is detected, McByte’s relative performance gain over the baseline is slightly reduced when using public detections. However, even in this scenario, the mask-based association cue helps resolve ambiguities and improves overall tracking performance, as shown in Tab. 17.
Method HOTA MOTA IDF1
@@
baseline 49.0 44.5 55.3
McByte 50.1 44.6 56.0
Table 17: ByteTrack and McByte with private and public detections on MOT17-val
Appendix F McByte components more in detail
@@
In order to ensure reliable use of the mask, we apply several conditions, which we mention in Sec. 3.3 in the main paper. We describe them more in detail in the following subsections.
F.1 Tracklet mask visibility at the scene
@@
In some cases, a tracklet’s mask may not appear in the current frame. This can happen if the subject is completely or mostly occluded, or if the mask temporal propagator fails to predict the mask due to challenges like small visible areas, dense crowds, or poor lighting conditions. When no mask is predicted for a tracklet, its associations with detections are calculated using intersection over union (IoU), as done in the baseline [38].
F.2 Mask confidence

### v3

@@ Appendix B State-of-the-art comparison with transformer-based and other types of method
MOTR [38] 57.8 68.6 73.4
MeMOTR [13] 58.8 71.5 72.8
MOTRv2 [41] 62.0 75.0 78.6
MOTIP [14] 59.2 71.2 75.5
@@
SUSHI [5] 66.5 83.1 81.1
Joint detection and tracking
FairMOT [39] 59.3 72.3 73.7
RelationTrack [37] 61.0 75.8 75.6
CenterTrack [42] 52.2 64.7 67.8
@@
with parameter tuning per sequence
ByteTrack [40] 63.1 77.3 80.3
StrongSORT++ [12] 64.4 79.5 79.6
OC-SORT [4] 63.2 77.5 78.0
Deep OC-SORT [27] 64.9 80.6 79.4
Hybrid-SORT [36] 64.0 78.7 79.9
Tracking-by-detection
without parameter tuning per sequence
ByteTrack [5] 62.8 77.1 78.9
C-BIoU [35] * 62.4 77.1 79.5
McByte (ours) 64.2 79.4 80.2
@@
Method HOTA IDF1 MOTA
@@
MOTR [38] 54.2 51.5 79.7
MeMOTR [13] 63.4 65.5 85.4
MOTRv2 [41] 73.4 76.0 92.1
MOTIP [14] 67.5 72.2 90.3
@@
SUSHI [5] 63.3 63.4 88.7
Joint detection and tracking
FairMOT [39] 39.7 40.8 82.2
CenterTrack [42] 41.8 35.7 86.8
@@
ByteTrack [40] 47.7 53.9 89.6
OC-SORT [4] 55.1 54.9 92.2
Deep OC-SORT [27] 61.3 61.5 92.3
C-BIoU [35] * 45.8 52.0 88.4
StrongSORT++ [12] 55.6 55.2 91.1
Hybrid-SORT [36] 65.7 67.4 91.8
McByte (ours) 67.1 68.1 92.9
@@
There exist MOT methods outside the tracking-by-detection domain manifesting performance differences, but usually these methods are not directly comparable, because they make certain hypotheses, e.g. global optimization on the whole video. At the same time, these methods might perform visibly worse on some benchmarks as we discuss below. On the contrary, we stress that McByte performs well on all the discussed benchmarks (Secs. 4.3 and 4.4 of the main paper). McByte is a tracking-by-detection approach, which is the main focus of our work. For an additional reference, though, we also list performance of the transformer-based, global optimization, and joint detection and tracking methods.
@@
Tab. 10 shows extended state-of-the art comparison on MOT17 [28] test set. Note that analogously to the main paper, we also put the result of ByteTrack [40] not being tuned per sequence as reported in [5]. Transformer-based methods perform visibly lower than the tracking-by-detection methods (including ours) as they struggle with the subjects frequently entering and leaving the scene. In contrast, SUSHI [5], which is a powerful global optimization approach, reaches highly satisfying performance. However, it accesses all the video frames at same time while processing detections and associating the tracklets, which makes it impossible to run in online settings. This makes the performance not directly comparable and therefore, we do not include it in the main paper. Current state-of-the-art joint detection and tracking methods generally perform lower than the tracking-by-detection methods. In this paradigm, the detection and association step is performed jointly. In our method, we perform these two steps separately and focus on the association part.
@@
Tab. 11 presents extended state-of-the-art comparison on DanceTrack [33] test set. As in this dataset the subjects remain mostly at the scene, the transformer-based methods performance is more satisfying. The performance of transformer-based methods can be both higher [41, 14] or lower [38, 13] compared to the the tracking-by-detection methods. For similar reasons, the global optimization method, SUSHI [5] can also perform higher than the other tracking-by-detection methods on this dataset, or lower, e.g. when compared to our method. On this dataset, joint detection and tracking methods also manifest lower performance than the tracking-by-detection methods. As in the case of the MOT17 benchmark, the performance of the mentioned different method types is not directly comparable and thus not included in the main paper.
Appendix C More information on C-BIoU
C.1 Our C-BIoU implementation and its performance
The C-BIoU [35] paper does not provide a public implementation or sufficient details for code reproduction. We implement C-BIoU based on the available information and report the best results on the datasets evaluated by the original algorithm. These results, marked with an asterisk, are shown in the main paper (Tabs. 3, 4 and 5) and in Tabs. 10, 11 and 12 of this supplementary material.
Based on the descriptions in the C-BIoU work [35], especially regarding tracklet management and tracklet-detection association, the core of the method strongly resembles ByteTrack [40]. Thus, we extend ByteTrack by incorporating cascaded buffered intersection over union [35] for the association process. Where C-BIoU does not specify parameters, we default to those used in ByteTrack. For the MOT17 test set, we do not tune detection confidence parameters per sequence (referred in Sec. 2.2 of the main paper) and apply a fixed 0.6 threshold for all sequences and datasets. This is the default value of ByteTrack, i.e. when ByteTrack is run on datasets without tuning per sequence. Regarding the two buffering scales in C-BIoU, we use the values 0.3 and 0.4, which were reported to result in the best HOTA on the DanceTrack validation set in the C-BIoU paper [35]. Although we also test the values 0.3 and 0.5 (as mentioned in the framework figure from C-BIoU), the 0.3 and 0.4 combination performs slightly better in our experiments.
Additionally, following C-BIoU’s supplementary material, we adopt the most optimal max_age values per dataset. The max_age parameter defines the maximum lifespan (in frames) of a tracklet that has not been matched to a detection before it is terminated and removed. For DanceTrack [33], we set max_age to 100, and for SoccerNet-tracking 2022 [9], we set it to 60. In case of MOT17 [28], where this value is not directly specified, we test three different max_age values: 20, 60, and 100 (as listed in C-BIoU’s supplementary material). We select the value that results in the highest performance on the MOT17 validation set, which is 60, and then use this value for testing on the MOT17 test set to report the final results.
The results are presented in Tabs. 3, 4 and 5 of the main paper. Compared to the baseline, ByteTrack [40], the improvement is modest (Tab. 5) or shows slight degradation (Tabs. 3 and 4). This could be due to tuning issues, as many parameters were not specified in the C-BIoU paper. Specifically, with buffered intersection over union (IoU), the association search space expands, leading to more potential associations and lower IoU-based distances. However, the matching thresholds for the linear association steps (using the Hungarian Algorithm [21]), which filter out unlikely matches, are not provided. Default thresholds from ByteTrack [40] may allow more matches, introducing unresolved ambiguities and potentially degrading performance.
C.2 C-BIoU with temporally propagated mask as an association cue
Method HOTA IDF1 MOTA
DanceTrack test set
C-BIoU [35] * 45.8 52.0 88.4
C-BIoU [35] * with our TP mask cue 67.6 70.2 92.4
SoccerNet-tracking 2022 test set
C-BIoU [35] * 72.7 76.4 95.4
C-BIoU [35] * with our TP mask cue 84.6 79.0 98.2
Table 12: Comaprison of another tracking algorithm [35], with and without our approach of temporally propagated (TP) mask as an association cue (Sec. C.2).
We also evaluate temporally propagated (TP) mask as an association cue within the C-BIoU [35] tracker to demonstrate the generalizability and advantage of our approach within MOT. We follow exactly the same steps as with incorporating the TP mask in our baseline [40] for McByte.  Tab. 12 includes the performance comparison of the C-BIoU tracker with and without the TP mask as a cue. The results show that our TP mask-based approach is beneficial beyond single baseline and that it can improve MOT performance of other trackers as well.
As in the other result tables, the asterisk means that no official implementation is published and that we reproduce the C-BIoU method based on the available information, described in detail in Sec. C.1. MOT17 [28] and KITTI-tracking [16] test evaluation servers allow limited submissions which we use for McByte and our baseline [40]. Therefore, we report the results of this study on DanceTrack [33] and SoccerNet-tracking 2022 [9] test sets.
@@ C.2 C-BIoU with temporally propagated mask as an association cue
Figure 5: An example of detection quality difference. It can be seen that considerably more bounding boxes are missing in case of public detections, which negatively impacts the MOT performance. Input image data from [28], sequence MOT17-04, last frame (1050). Best seen in color.
Appendix D McByte and baseline on public detections
@@
We also evaluate McByte and the baseline [40] on the MOT17 [28] validation set using the public detections provided by the dataset authors. The results are shown in Tab. 13, using the FRCNN model variant [30]. There are significant performance differences for both the baseline and McByte when using public detections, which are of lower quality compared to the private YOLOX [15] detections pre-trained on MOT17 by the baseline [40]. Public detections tend to miss more objects, leading to many false negatives.
@@
An example comparing the quality of private and public detections is illustrated in Fig. 5. Since tracklets and their temporally propagated masks can only be initiated when a bounding box is detected, McByte’s relative performance gain over the baseline is slightly reduced when using public detections. However, even in this scenario, the temporally propagated mask-based association cue helps resolve ambiguities and improves overall tracking performance, as shown in Tab. 13.
Appendix E McByte components more in detail
Method HOTA IDF1 MOTA
@@ Appendix E McByte components more in detail
Baseline 49.0 55.3 44.5
McByte 50.1 56.0 44.6
Table 13: McByte and baseline [40] with private and public detections on MOT17-val
E.1 Mask processing and management
We provide more details on mask handling introduced in Sec. 3.2 of the main paper. New objects might appear in any frame on the scene. It involves initiating new tracklets and necessitates creating new segmentation masks. We use an image segmentation model and provide the bounding boxes of new tracklets and the currently processed frame to create new masks. Further, a mask temporal propagator receives new initial masks and the frame at which they were created. Subsequently, the propagator receives the next frame, for which it returns the updated mask predictions, e.g. based on the object movement and scenery change. Each tracklet has an ID and each mask included within the mask temporal propagator also gets its ID. We link masks to the tracklets based on their IDs.
Tracklets might also be marked inactive and hence removed at any frame, e.g. when a tracklet has not been matched to any detection for a long time. It induces removing the tracklets and necessitates purging the masks from the mask temporal propagator. As masks are removed from the propagator, the IDs of the remaining masks change (shift) due to the memory allocation mechanisms [6, 8]. We update the links between tracklet IDs and mask IDs to consistently keep the same masks assigned to the tracklets. Tracklet IDs are immutable. For our experimentation, we use SAM [20] as the image segmentation model and Cutie [8] as the mask temporal propagator.
The mask handling process is illustrated in Fig. 6. We process each frame of the video sequence separately. We check if there are any new initialized tracklets from the previously processed frame. If they are strongly occluded, their IDs are added to the waiting list. The tracklet positions change over the frames and they might become less occluded during the next frames. We take currently non-occluded tracklets, both newly initialized and from the waiting list, and create their masks with the image segmentation model. We include the new masks within the mask temporal propagator.
Further, we check if any tracklets have been removed at the previous frame and if so, we purge their masks from the mask temporal propagator. Regardless of adding or removing any masks, we perform mask temporal propagation to get the updated mask predictions for the current frame. The temporally propagated masks are passed to the tracker where they are used to improve the tracklet-detection association process. The process of temporally propagated masks influencing the association process is illustrated in Fig. 2 (main paper) and described in Sec. 3.3 (main paper). After the association, the tracker returns updated tracklets. Output tracklets and masks from the current frame processing constitute the input for the next frame processing. During the whole process, state of the tracklets influences the mask management, i.e. the mask creation and purging. At the same time, the masks influence the tracklet association and updating, improving the MOT performance.
Figure 6: Diagram illustrating the mask processing and management process as described in Sec. E.1. Tracklets influence the mask management process, while masks influence the tracklet association and updating. The process of temporally propagated masks influencing the association process (the blue part) is illustrated in Fig. 2 (main paper).
E.2 Tracklet mask visibility at the scene
@@
In order to ensure reliable use of the temporally propagated (TP) mask, we apply several conditions, which we mention in Sec. 3.3 in the main paper. We describe them more in detail in the following subsections starting with tracklet mask visibility at the scene.
@@
In some cases, a tracklet’s TP mask may not appear in the current frame. This can happen if the subject is completely or mostly occluded, or if the mask temporal propagator fails to predict the mask due to challenges like small visible areas, dense crowds, or poor lighting conditions. When no mask is predicted for a tracklet, its associations with detections are calculated using intersection over union (IoU), as done in the baseline [40].
E.3 Mask confidence